OpenStudy (nerdgamer_):

Medal to whoever helps and a fan. Which improper fraction is equivalent to two and three over nine? fourteen over nine twenty-nine over nine twenty-one over nine fifteen over nine

OpenStudy (nerdgamer_):

I know its easy but i dont understand.

jigglypuff314 (jigglypuff314):

to convert 2 3/9 into an improper fraction, know that 2 = 18/9 right? because 18 divided by 9 = 2 so then what is 18/9 + 3/9 \[\frac{ 18 }{ 9 } + \frac{ 3 }{ 9 } =\]

OpenStudy (nerdgamer_):

right so it would be

OpenStudy (nerdgamer_):

hold on

OpenStudy (nerdgamer_):

I still need the answer i dot understant

OpenStudy (nerdgamer_):

hello are you there?

OpenStudy (nerdgamer_):

=what?

jigglypuff314 (jigglypuff314):

yeah, okay, to add fractions, the denominators have to be the same. (denominator is the number below the fraction line) in this case, they are both 9 then you add the numerators (the numbers above the fraction line) so what is 18 + 3 = ?

OpenStudy (nerdgamer_):

21

OpenStudy (nerdgamer_):

i know what a denominater is lol.

jigglypuff314 (jigglypuff314):

yes, so since 18+3=21 then 18/9 + 3/9 = 21/9

OpenStudy (nerdgamer_):

ohhh ok thanks soo much i give you a medal and fan ok.

jigglypuff314 (jigglypuff314):

you're welcome and thanks ;)

OpenStudy (nerdgamer_):

ok

OpenStudy (nerdgamer_):

your welcom